OpenStudy (anonymous):

2sin^2theta-sin theta= 0 with intervals 0 less than or equal to theta less than or equal to 360

OpenStudy (zehanz):

\(2\sin^2 \theta - \sin \theta=0 \Leftrightarrow \sin \theta(2\sin \theta - 1)=0\) Now solve separately sinθ = 0 and 2sinθ−1=0.

OpenStudy (anonymous):

Sin theta= 0 and sin=1/2

OpenStudy (zehanz):

Yes, so what can theta be then?